OpenStudy (anonymous):

What is the value of x? -15-x=-6

Mehek (mehek14):

add -6 + 15

OpenStudy (anonymous):

9

Mehek (mehek14):

so that's the answer

OpenStudy (anonymous):

=> -15-x=-6 => therefore, -x=-6+15 => -x=9 => therefore, x=-9

OpenStudy (anonymous):

9 or negative 9?

OpenStudy (anonymous):

9 is negative:)

Mehek (mehek14):

positive

Mehek (mehek14):

greater number = +15 so the answer is positive

pooja195 (pooja195):

Negative.

Mehek (mehek14):

oh wait nvm -x >_>

Mehek (mehek14):

im dun

OpenStudy (anonymous):

@mehek 15 is negative too

OpenStudy (anonymous):

Its negative @Mehek14 see from my method

Mehek (mehek14):

ik but when u add it's9 but i didn't see -x = 9

OpenStudy (alekos):

how did you get a score of 99?

OpenStudy (anonymous):

Thank you all! Off to shave my face now!

Mehek (mehek14):

XD

pooja195 (pooja195):

@alekos Be nice.

OpenStudy (anonymous):

@alekos no one said 99

OpenStudy (anonymous):

Yea :) So now the minus sign of x transposes to the other side (LHS=9) and thus becomes x=-9

Mehek (mehek14):

@alekos people make mistakes

OpenStudy (alekos):

thats not what i meant

Mehek (mehek14):

@vishnupriya7 thank you :)

OpenStudy (anonymous):

What is the value of x? -5x=-35

OpenStudy (alekos):

OK, fair enough

OpenStudy (anonymous):

Its okay everybody, yea everybody make mistake and everybody improves :)

Mehek (mehek14):

-35/-5

pooja195 (pooja195):

divide both sides by -5

OpenStudy (anonymous):

7?

pooja195 (pooja195):

Well done

OpenStudy (anonymous):

thanks

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

What is the value of x? 4x-9=27

pooja195 (pooja195):

add 9 to both sides 27+9=?

OpenStudy (anonymous):

36

pooja195 (pooja195):

good 36/4=?

OpenStudy (anonymous):

9

pooja195 (pooja195):

Good so x=9

OpenStudy (anonymous):

Awesome. I'm gonna close this question and open a new one

pooja195 (pooja195):

ok ^-^

hartnn (hartnn):

Durga mata ki jai!